Novak Company’s net income for 2017 is $53,400. The only potentially dilutive securities outstanding were 1,100 options issued during 2016, each exercisable for one share at $6. None has been exercised, and 10,300 shares of common were outstanding during 2017. The average market price of Novak’s stock during 2017 was $20.

(a) Compute diluted earnings per share. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share $


(b) Assume the same facts as those assumed for part (a), except that the 1,100 options were issued on October 1, 2017 (rather than in 2016). The average market price during the last 3 months of 2017 was $20. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share
Answer #1

Answer (a):- Calculation of diluted Earning per share(EPS)

No. of potential shares in options = 1100 - (1100*6)/20

= 770 Shares

So, total shares = 10,300+770= 11,070

Diluted EPS = 53,400/11,070 = 4.82

Answer (b):-If issued on Oct, 1 2017

= 770*3/12

= 192.50

So, total shares = 10,300+192.50= 10,492.50

Diluted EPS = 53,400/10,492.50 = 5.09